### v2

# Large Multi-modality Model Assisted AI-Generated Image Quality Assessment

Abstract.
11footnotetext: Corresponding author.††footnotetext: All of the authors are with Department of Electronic Engineering, Shanghai Jiao Tong University, Shanghai 200240 China. Puyi Wang is also with the Zhiyuan College, Shanghai Jiao Tong University.Traditional deep neural network (DNN)-based image quality assessment (IQA) models leverage convolutional neural networks (CNN) or Transformer to learn the quality-aware feature representation, achieving commendable performance on natural scene images. However, when applied to AI-Generated images (AGIs), these DNN-based IQA models exhibit subpar performance. This situation is largely due to the semantic inaccuracies inherent in certain AGIs caused by uncontrollable nature of the generation process. Thus, the capability to discern semantic content becomes crucial for assessing the quality of AGIs. Traditional DNN-based IQA models, constrained by limited parameter complexity and training data, struggle to capture complex fine-grained semantic features, making it challenging to grasp the existence and coherence of semantic content of the entire image. To address the shortfall in semantic content perception of current IQA models, we introduce a large Multi-modality model Assisted AI-Generated Image Quality Assessment (MA-AGIQA) model, which utilizes semantically informed guidance to sense semantic information and extract semantic vectors through carefully designed text prompts. Moreover, it employs a mixture of experts (MoE) structure to dynamically integrate the semantic information with the quality-aware features extracted by traditional DNN-based IQA models. Comprehensive experiments conducted on two AI-generated content datasets and two traditional IQA datasets show that MA-AGIQA achieves state-of-the-art performance, and demonstrate its superior generalization capabilities on assessing the quality of AGIs. The code is available at https://github.com/wangpuyi/MA-AGIQA.
1. Introduction
The rapid advancement of artificial intelligence (AI) has led to a proliferation of AI-generated images (AGIs) on the Internet. However, current AI-driven image generation systems often produce multiple images, necessitating manual selection by users to identify the best ones. This labor-intensive process is not only time-consuming but also a significant barrier to fully automating image processing pipelines. Visual quality, as an important factor to select attractive AGIs, has gained lots of attention in recent years (Li et al., 2023b, 2024a). In this paper, we focus on how to evaluate the visual quality of AGIs, which on the one hand can be used to filter high-quality images from generation systems and on the other hand, can sever as reward function to optimize image generation models (Black et al., 2023), propelling progress in the field of AI-based image generation techniques.
While a substantial number of deep neural network (DNN)-based image quality assessment (IQA) models, such as HyperIQA (Su et al., 2020), MANIQA (Yang et al., 2022), DBCNN (Zhang et al., 2018), etc., have been developed, these models were specifically designed for and trained on natural scene images. When applied directly to AGIs, these models often exhibit poor performance. This is due to the fact that quality assessment of natural images primarily targets issues such as blur, noise, and other forms of degradation caused by photography equipment or techniques, which are not applicable to AGIs as they do not undergo such degradation during the generation process. Therefore, overemphasizing factors like blur or noise during the evaluation of AGIs is inappropriate.
As shown in Figure 1, AI-generated images, derived from advanced image generative models such as generative adversarial networks (GANs) (Ma et al., 2023), diffision (Ho et al., 2020) and related variant (Rombach et al., 2022b; Xu et al., 2018; Ramesh et al., 2021, 2022; dreamlike art, 2023; DeepFloyd, 2023; Holz, 2023; PlaygroundAI, 2023), often exhibit issues not commonly found in naturally captured images. Visual quality of AGIs depends not only on basic visual features such as noise, blur (Su et al., 2021; Li et al., 2022; Zhang et al., 2024), etc., but also on more intricate semantic perception (Li et al., 2024a), such as existence of reasonable semantic content, scene plausibility, and the coherence among objects (Wu et al., 2023b; Zhang et al., 2023; Li et al., 2024b; Wu et al., 2023c, 2024b). Although re-training existing IQA models on AGIs datasets leads to improved outcomes, it fails to achieve optimal performance. One reason is that traditional DNN models, especially early convolutional neural networks (CNNs), despite their notable achievements in tasks like image recognition and classification (Simonyan and Zisserman, 2015; He et al., 2015; Szegedy et al., 2014), still struggle to grasp the fine-grained semantic content of images (Zhang et al., 2021). What’s more, traditional DNN based IQA models fail to capture the intrinsic characteristics essential for assessing image quality and thus exhibit poor generalization abilities. Hence, we argue that the quality assessment models of AGIs are still in their infancy and need further exploration.
To address the issue of semantic awareness, we resort to large multi-modality models (LMMs). Because LMMs is typically pre-trained on large-scale datasets and has already learned a rich set of joint visual and language knowledge, it can effectively capture the fine-grained semantic features relevant to input prompts. However, LMMs perform excellently in high-level visual understanding tasks (Achiam et al., 2023; Li et al., 2023a), yet they do not perform well on tasks that are relatively simple for humans, such as identifying structural and textural distortions, color differences, and geometric transformations (Wu et al., 2024a). In contrast, traditional deep learning networks excel at perceiving low-dimensional features and can fit better to the data distribution of specific task (Hornik et al., 1989). Therefore, the idea of combining LMMs with traditional deep learning networks is a natural progression.
In this paper, we introduce a large Multi-modality model Assisted AI-Generated Image Quality Assessment (MA-AGIQA) framework, which enhances the capacity of traditional DNN-based IQA models to understand semantic content by incorporating LMM. Our approach initially repurposes a DNN, MANIQA (Yang et al., 2022), as an extractor for quality-aware features and establishes it as the training backbone for the MA-AGIQA framework. Subsequently, we guide a LMM, mPLUG-Owl2 (Ye et al., 2023), to focus on fine-grained semantic information through meticulously crafted prompts. We then extract and store the last-layer hidden vector from mPLUG-Owl2, merging it with features extracted by MANIQA to infuse the model with rich semantic insights. Finally, we employ a MoE to dynamically integrate quality-aware features with fine-grained semantic features, catering to the unique focal points of different images.
As an evidence that our MA-AGIQA model are clearly superior to traditional DNN-based methods when it comes to AGIs, a subset with grainy images (the concept of ”grainy” is depicted like the grainy orange cat in Figure 1) is sampled out from test set of AIGCQA-20k. In this subset, MANIQA achieves an SRCC of 0.2545, which is 70.0% lower than its overall SRCC performance (0.8507). In contrast, our model achieves an SRCC of 0.8364 in this subset and has an overall performance of 0.8644. It demonstrates that our model possesses a significantly enhanced understanding of AGIs, particularly those whose quality is deeply intertwined with semantic elements. Besides, MA-AGIQA achieves SRCC values of 0.8939 and 0.8644 on the AGIQA-3k and AIGCQA-20k datasets, respectively, exceeding the state-of-the-art models by 2.03% and 1.37%, and also demonstrates superior cross-dataset performance.
Our contributions are three-fold:
We systematically analyze the issue of traditional DNN-based IQA lacking the ability to understand the semantic content of AGIs, emphasizing the importance of incorporating semantic information into traditional DNN-based IQA models.
We introduce the MA-AGIQA model, which incorporates LMM to extract fine-grained semantic features and dynamically integrates these features with traditional DNN-based IQA models.
We evaluate the MA-AGIQA model on two AI-generated IQA datasets and two traditional IQA datasets. Experimental results demonstrate that our model surpasses current state-of-the-art methods without extra training data and also showcases superior cross-dataset performance. Extensive ablation studies further validate the effectiveness of each component.
2. Related Work
Traditional IQA models. In the field of No-Reference Image Quality Assessment (NR-IQA) (Zhai and Min, 2020), traditional models primarily fall into two categories: handcrafted feature-based and DNN-based.
Models based on handcrafted features, such as BRISQUE (Mittal et al., 2012a), ILNIQE (Zhang et al., 2015), and NIQE (Mittal et al., 2012b), primarily utilize natural scene statistics (NSS) (Mittal et al., 2012a, b) derived from natural images. These models are adept at detecting domain variations introduced by synthetic distortions, including spatial (Mittal et al., 2012a, b; Zhang et al., 2015), gradient (Mittal et al., 2012a), discrete cosine transform (DCT) (Saad et al., 2012), and wavelet-based distortions (Moorthy and Bovik, 2011). However, despite their effectiveness on datasets with type-specific distortions, these handcrafted feature-based approaches exhibit limited capabilities in modeling real-world distortions.
With the advent of deep learning, CNNs have revolutionized many tasks in computer vision. (Kang et al., 2014) is pioneer in applying deep convolutional neural networks to NR-IQA. Its methodology employs CNNs to directly learn representations of image quality from raw image patches, bypassing the need for handcrafted features or a reference image. Following this, DBCNN (Zhang et al., 2018) introduces a deep bilinear CNN for NR-IQA (Zhai and Min, 2020), innovatively merging two CNN streams to address both synthetic and authentic image distortions separately. Furthermore, HyperIQA (Su et al., 2020), a self-adaptive hyper network, evaluates the quality of authentically distorted images through a novel three-stage process: content understanding, perception rule learning, and quality prediction.
The success of Vision Transformers (ViT) (Dosovitskiy et al., 2020) in various computer vision tasks has led to significant advancements. In the realm of IQA, IQT (You and Korhonen, 2021) leverages the combination of reference and distorted image features, extracted by CNNs, as inputs for a Transformer-based quality prediction task. MUSIQ (Ke et al., 2021) utilizes a Transformer to encode distortion image features across three scales, addressing the challenge of varying input image sizes during training and testing. TReS (Golestaneh et al., 2022) introduces relative ranking and self-consistency loss to capitalize on the abundant self-supervisory information available, aiming to decrease the network’s sensitivity. What’s more, MANIQA (Yang et al., 2022) explored multi-dimensional feature interaction, utilizing spatial and channel structural information to calculate a non-local representation of the image, enhancing the model’s ability to assess image quality comprehensively.
LMMs for IQA. Recent methodologies employing LMMs for IQA either utilize LMMs in isolation or combine them with DNNs as feature extractors to enhance performance. (Qu et al., 2024) introduces an innovative image-prompt fusion module, along with a specially designed quality assessment token, aiming to learn comprehensive representations for AGIs, providing insights from image-prompt alignment. However, the evaluation of AGIs in practical scenarios often does not involve prompts and image-prompt alignment is more significant for assessing the capabilities of generative models rather than images quality. CLIPIQA (Wang et al., 2023) signifies a breakthrough in assessing image quality and perception by harnessing the strengths of CLIP (Radford et al., 2021) models. This method bridges the divide between measurable image quality attributes and subjective perceptions of quality without necessitating extensive labeling efforts. Nonetheless, their (Qu et al., 2024; Wang et al., 2023) dependence on visual-text similarity for quality score prediction often constrains their performance, rendering it marginally less effective compared to methods that exclusively focus on visual analysis. What’s more, Q-Bench (Wu et al., 2023b) innovates with a softmax strategy, allowing LMMs to deduce quantifiable quality scores. This is achieved by extracting results from softmax pooling on logits corresponding to five quality-related tokens. And Q-Align (Wu et al., 2023a) employs strategic alignment techniques to foster accuracy. Expanding further, (Wu et al., 2024a) delves into enhancing the assessment of AGIs by focusing on optimizing individual text prompts to leverage the intrinsic capabilities of LMMs, aiming to provide a more nuanced understanding and evaluation of image quality of AGIs. However, these methods, while notable, fall short of achieving satisfying efficacy, leaving considerable room for improvement.
3. Method
As depicted in Figure 2, framework of MA-AGIQA is structured into three sections. Section 3.1 introduces our adoption of a DNN, specifically MANIQA (Yang et al., 2022), tailored for the AGIs quality assessment task, serving as our primary training backbone. In Section 3.2, we incorporate the LMM mPLUG-Owl2 (Ye et al., 2023) as a feature extractor. This component is crucial for acquiring fine-grained semantic features via carefully crafted text prompts. Lastly, Section 3.3 utilizes a MoE structure for adaptive feature fusion, which addresses the variability in focal points across differenct images and ensures that the most salient features are emphasized. Further details are elaborated below.
3.1. Quality-aware Feature Extraction
To leverage the capability of DNNs to adapt to the data distribution of specific tasks, we employ MANIQA (Yang et al., 2022) as a quality-aware feature extractor. MANIQA enhances the image quality evaluation by applying attention mechanisms across both the channel and spatial dimensions, thereby increasing the interaction among various regions of the image, both globally and locally. This approach generates projections and () for a given image, and the final rating of the whole image is determined through the sum of multiplication of by , which can be illustrated as Equation (1):
rating |
where and are one dimensional vectors.
However, directly applying MANIQA to the quality assessment of AGIs presents challenges, as illustrated in Figure 3. Image (a) displays a complex, symmetrical pattern, devoid of meaningful semantic content. Image (b) features incoherent areas, such as two grey holes in the sky that are inconsistent with the common sense. The blurriness and fuzziness of the man’s face in image (c) along the edges significantly impair human perception. Conversely, image (d), despite its severe graininess, retains its semantic integrity, representing an appealing artistic form. Traditional DNN-based models like MANIQA, lacking the capacity to comprehend semantic content, tend to overestimate the quality of images (a), (b), and (c), resulting in scores much higher than the ground truth. However, these images should be rated as low quality due to the poor viewing experience they offer. For image (d), traditional DNN-based models focus excessively on the graininess, mistaking it for a flaw, and assign a score significantly lower than the ground truth. This highlights the critical need for incorporating semantic information into the quality assessment of AGIs by traditional DNN-based models.
To address this issue, modifications were made so that the generated and no longer produce a rating. Instead, they yield a quality-aware feature , setting the stage for the subsequent fusion with features extracted by LMM. is generated as:
During the training phase, the parameters of modified MANIQA are continuously updated. This refinement process ensures that MANIQA can extract features more relevant to the quality of AGIs and facilitates a more seamless integration between MANIQA and LMM, leading to superior outcomes.
3.2. Fine-grained Semantic Feature Extraction
LMMs are utilized to assess whether different parts of an image form a cohesive whole and whether the elements within the picture are semantically coherent (Gao et al., 2023; OpenAI, 2023; Liu et al., 2023). mPLUG-Owl2 (Ye et al., 2023) employs a modality-adaptive language decoder to handle different modalities within distinct modules, which mitigates the issue of modality interference. Given the importance of effectively guiding the model through textual prompts to elicit the desired output, we have selected mPLUG-Owl2 as our feature extractor.
We consider the application of mPLUG-Owl2 in the following aspects of semantic content:
Existence of Semantic Content. The importance of semantic content in an image lies in its ability to convey a clear and meaningful message to the viewer. An image lacking in semantic content may be difficult to understand, fail to effectively convey its intended message, reducing audience engagement and satisfaction.
Coherence of Semantic Content. The coherence of semantic content in an image relates to whether the generated image can provide a coherent, logically sound visual experience for human viewers. When the various parts of an image are semantically consistent, it is better able to convey a clear story, emotion, or message. In contrast, any inconsistency in the primary focus of images will greatly detract from their quality and convey a significantly negative impression.
Consequently, we try to propose the rational design of prompts leading LMMs to obtain those image semantic content. mPLUG-Owl2 possess the ability to understand fine-grained semantic contents, but without carefully designed input prompts, some prompts, such as ”Please evaluate if the image quality is compromised due to violations of common human sense or logic?” although expresses the desire for the model to assess whether the semantic content of the image contradicts human perception, would lead to unsatisfactory results. So we meticulously designed two prompts to guide the LMM, denoted as and respctively. Specifically,
”Evaluate the input image to determine if its quality is compromised due to a lack of meaningful semantic content.”
”Evaluate if the image quality is compromised due to violations of coherence.”
corresponding to the existence of semantic content and the coherence of semantic content in images. Test results, as shown in Figure 4 using the mPLUG-Owl2 official demo111https://modelscope.cn/studios/iic/mPLUG-Owl2/summary, have proven these questions to be effective.
However, the text format is not conducive to being utilized by MANIQA to impart semantic insights. To bridge this gap, it’s essential to transform the information into a format that MANIQA can easily leverage. So we extract features from the final layer of mPLUG-Owl2’s hidden layers, achieving an accessible embedded representation of the LMM’s output. This output is a tensor with dimensions of [token_length, hidden_size], where ”token_length” represents the number of output tokens, and ”hidden_size” denotes the dimensionality of the hidden layer representations associated with each token. For mPLUG-Owl2, the hidden_size is set to 4096. Subsequently, we conduct an averaging operation across the token dimension, yielding a vector with dimensions 1x4096. This vector then serves as the basis for further feature fusion procedures. The process can be represented as Equation (3) :
where represents a hidden vector of token corresponding to , and denotes mPLUG-Owl2.
It is important to note that throughout the entire training and testing process, the parameters of mPLUG-Owl2 are fixed. Because mPLUG-Owl2 is typically pre-trained on large-scale datasets and has already learned a rich set of joint visual and language knowledge, it can effectively capture the fine-grained semantic information relevant to input prompts, even with fixed parameters. Additionally, fine-tuning LMMs in every training iteration would significantly increase training time. Using it solely as a feature extractor significantly reduces computational costs, making the training process more efficient. So, we pre-obtain and save the semantic content features of each image in advance.
3.3. Adaptive Fusion Module
Given the complex influence of color, composition, details, semantic content, and other factors on image quality, simply concatenating the extracted features may not always yield the best results. To dynamically fuse a variety of complementary features, we propose the adaptive fusion module (AFM) for organic feature integration. This process can be divided into two main parts. The first part involves transforming the extracted features into a unified vector space of the same dimension, allowing for vector fusion operations. Specifically, for quality-aware features, this transformation block simply applies a Fc layer, transforming them to the same dimension as the original features (1x784). For fine-grained semantic features, it uses a Fc layer to project them onto a 1x784 dimension, followed by a relu activation layer and a dropout layer to enhance the network’s expressive power and generalization. The second part employs a MoE to dynamically fuse the three features. The MoE’s gating network takes the transformed three features as input and outputs dynamic weights , corresponding to three features’ contributions to image quality. Structurally, this gating network comprises a Fc layer and a sigmoid layer. The final image quality representation vector can be obtained through a weighted sum of the three feature vectors. Following the denotation which sign the three features as , , , this process can be represented as:
where is the transformation block of feature , and is the gating network’s mapping function, is the original extracted feature and is the transformed feature. is the dimension space of . Finally, we obtain the final image quality score output through a simple regression layer, consisting of a Fc layer.
Type | Method | AGIQA-3k | AIGCQA-20K | KonIQ-10k | SPAQ | ||||
SRCC | PLCC | SRCC | PLCC | SRCC | PLCC | SRCC | PLCC | ||
Handcrafted feature-based | BRISQUE (Mittal et al., 2012a) | 0.4726 | 0.5612 | 0.1663 | 0.3580 | 0.7050 | 0.7070 | 0.8021 | 0.8056 |
NIQE (Mittal et al., 2012b) | 0.5236 | 0.5668 | 0.2085 | 0.3378 | 0.5512 | 0.4881 | 0.7030 | 0.6700 | |
ILNIQE (Zhang et al., 2015) | 0.6097 | 0.6551 | 0.3359 | 0.4551 | 0.4531 | 0.4676 | 0.7194 | 0.6540 | |
LMM-based | CLIPIQA (Wang et al., 2023) | 0.6524 | 0.6968 | 0.4147 | 0.6459 | 0.6957 | 0.7270 | 0.7385 | 0.7354 |
CLIPIQA+ (Wang et al., 2023) | 0.6933 | 0.7493 | 0.4553 | 0.6682 | 0.8756 | 0.8786 | 0.8590 | 0.8796 | |
Q-Align (Wu et al., 2023a) | 0.6728 | 0.6910 | 0.6743 | 0.6815 | 0.9223 | 0.9113 | 0.8875 | 0.8860 | |
Traditional DNN-based | HyperIQA (Su et al., 2020) | 0.8509 | 0.9049 | 0.8162 | 0.8329 | 0.9041 | 0.9152 | 0.9155 | 0.9188 |
MANIQA (Yang et al., 2022) | 0.8618 | 0.9115 | 0.8507 | 0.8870 | 0.9300 | 0.9460 | 0.9229 | 0.9278 | |
DBCNN (Zhang et al., 2018) | 0.8263 | 0.8900 | 0.8054 | 0.8483 | 0.8443 | 0.8624 | 0.9099 | 0.9273 | |
StairIQA (Sun et al., 2023) | 0.8343 | 0.8933 | 0.7899 | 0.8428 | 0.9209 | 0.9362 | 0.9238 | 0.9273 | |
BAID (Yi et al., 2023) | 0.1304 | 0.2030 | 0.1652 | 0.1483 | 0.2535 | 0.2658 | 0.3256 | 0.3232 | |
MUSIQ (Ke et al., 2021) | 0.8261 | 0.8657 | 0.8329 | 0.8646 | 0.8249 | 0.9376 | 0.8732 | 0.8688 | |
DNN with LMM | MA-AGIQA | 0.8939 | 0.9273 | 0.8644 | 0.9050 | 0.9339 | 0.9482 | 0.9278 | 0.9321 |
4. Experiments
4.1. Dataset and Evaluation Metrics
Dataset. Our model is evaluated on two AI-Generated image datasets (Li et al., 2023b, 2024a) and two traditional IQA datasets (Hosu et al., 2020; Fang et al., 2020). Note that AIGCQA-20k contains 20k images, but at the time of writing, only 14k images have been published and our experiments are conducted on these 14k images. Images in AIGCQA-20k are generated by 15 models, including DALLE2 (Ramesh et al., 2022), DALLE3 (Ramesh et al., 2022), Dream (dreamlike art, 2023), IF (DeepFloyd, 2023), LCM Pixart (Luo et al., 2023), LCM SD1.5 (Luo et al., 2023), LCM SDXL (Luo et al., 2023), Midjourney (Holz, 2023), Pixart (Chen et al., 2023), Playground (PlaygroundAI, 2023), SD1.4 (Rombach et al., 2022b), SD1.5 (Rombach et al., 2022b), SDXL (Rombach et al., 2022a) and SSD1B (Gupta et al., 2024). AGIQA-3k includes 2982 images. Images in AGIQA-3k are derived from six models, including GLIDE (Nichol et al., 2022), SD1.5, SDXL-2.2, Midjourney, AttnGAN (Xu et al., 2018), and DALLE2. During training, we split the entire dataset into 70 for training, 10 for validation, and 20 for testing. To ensure the same set of images in each subset when testing across different models, we set the same random seed during the split to ensure reproducibility.
Evaluation Metric. Spearman’s Rank-Order Correlation Coefficient (SRCC), Pearson’s Linear Correlation Coefficient (PLCC), the Kullback-Leibler Correlation Coefficient (KLCC), and the Root Mean Square Error (RMSE) are selected as metrics to measure monotonicity and accuracy. SRCC, PLCC, and KLCC range from 0 to 1.0, with larger values indicating better results. In our experiments, we employ the sum of SRCC and PLCC as the criterion for selecting the optimal validation case, and emphasize SRCC for comparing model performance.
4.2. Implementation Details
Our method is implemented based on PyTorch, and all experiments are conducted on 4 NVIDIA 3090 GPUs. For all datasets, we opt for handcrafted feature-based BRISQUE (Mittal et al., 2012a), NIQE (Mittal et al., 2012b) and ILNIQE (Zhang et al., 2015), DNN-based HyperIQA (Su et al., 2020), MANIQA (Yang et al., 2022), MUSIQ (Ke et al., 2021), DBCNN (Zhang et al., 2018), StairIQA (Sun et al., 2023), BAID (Yi et al., 2023), and LMM-based CLIPIQA (Wang et al., 2023), CLIPIQA+ (Wang et al., 2023) and Q-Align (Wu et al., 2023a). During the training process, all DNN-based models are trained for 30 epochs using MSE loss and validated after each training process. Handcrafted feature-based and LMM based models are used directly without training.
4.3. Comparison with SOTA methods
Table 1 lists the results of MA-AGIQA and 12 other models. On AGIs dataset, it has been observed that LMM-based models significantly outperform those that rely on handcrafted features. This superior performance is attributed to LMMs being trained on extensive datasets, which provides them with a robust understanding of images and enhances their generalizability. However, trained DNN-based models generally perform far better than the LMM-based models because DNN-based models tend to fit the data distribution of specific tasks better, thereby resulting in improved performance. Among these twelve models, the ViT-based MANIQA outperforms the other eleven models, and our method still significantly surpasses it on the same training and testing split with large margins. Despite being specifically designed for AGIs, MA-AGIQA outperforms all the compared methods on both NSIs datasets. This suggests that semantic information is also crucial for evaluating the quality of NSIs. It is worth noting that semantic information plays a more significant role in AGI quality assessment than in NSI quality assessment, as evidenced by the larger performance gap on AGIs datasets.
direction | SRCC | PLCC | KRCC | RMSE | |
HyperIQA | 20k3k | 0.6820 | 0.6806 | 0.4806 | 0.7352 |
3k20k | 0.6374 | 0.6547 | 0.4577 | 0.5414 | |
MANIQA | 20k3k | 0.7653 | 0.7980 | 0.5681 | 0.6041 |
3k20k | 0.7301 | 0.7975 | 0.5372 | 0.4402 | |
MA-AGIQA | 20k3k | 0.8053 | 0.8430 | 0.6083 | 0.5399 |
3k20k | 0.7722 | 0.8314 | 0.5777 | 0.4055 |
To evaluate the generalization capability of our MA-AGIQA model, we conducted cross-dataset evaluations. Table 2 shows that MA-AGIQA significantly outperforms the other two models, HyperIQA and MANIQA, which performed best on single datasets, with large margins. This superior performance can largely be attributed to the robust generalization capability of the LMM and the benefits of the MoE architecture, which excels in dynamically fusing features.
4.4. Ablation Study
MANIQA | Semantic Feature | Coherence Feature | SRCC | PLCC | KRCC |
✓ | 0.8800 | 0.9196 | 0.7031 | ||
✓ | 0.8662 | 0.9082 | 0.6823 | ||
✓ | 0.8661 | 0.9084 | 0.6821 | ||
✓ | ✓ | 0.8685 | 0.9108 | 0.6853 | |
✓ | ✓ | 0.8820 | 0.9197 | 0.7090 | |
✓ | ✓ | 0.8699 | 0.9102 | 0.6867 | |
✓ | ✓ | ✓ | 0.8939 | 0.9273 | 0.7211 |
Necessity of Fine-grained Semantic Features. To assess the benefits of integrating features extracted by mPLUG-Owl2 (Ye et al., 2023) into MANIQA (Yang et al., 2022), we carried out comprehensive ablation studies on each component and their various combinations, as detailed in Tables 3 and 4. Our findings indicate that using either the features extracted by the LMM alone or solely relying on a traditional network does not yield the best outcomes. In contrast, integrating one fine-grained semantic feature with the original MANIQA network can enhance the network’s performance. However, the optimal results were achieved by combining two features extracted by the LMM with MANIQA, which led to significant improvements on two AGIs datasets.
The marked enhancements achieved by incorporating two fine-grained semantic features suggest that LMM is adept at capturing nuanced, complex features that traditional models might overlook, fostering a more thorough understanding and assessment of AGIs quality. The results from these ablation experiments highlight the significant contribution of fine-grained semantic features.
Contribution of MoE. Table 5 demonstrates that incorporating the MoE structure, rather than simply concatenating three vectors, does improve network performance, albeit marginally. Specifically, on the AGIQA-3k dataset, we observed increases of 0.20, 0.17 in SRCC and PLCC, respectively. For the AIGCQA-20k dataset, the improvements were 0.67 and 0.95. The gains, although seemingly modest, highlight the potential of MoE structure in complex systems where integrating diverse expertise can yield better decision-making and predictive outcomes.
MANIQA | Semantic Feature | Coherence Feature | SRCC | PLCC | KRCC |
✓ | 0.8415 | 0.8877 | 0.6520 | ||
✓ | 0.8184 | 0.8345 | 0.6323 | ||
✓ | 0.8181 | 0.8343 | 0.6320 | ||
✓ | ✓ | 0.8540 | 0.8975 | 0.6671 | |
✓ | ✓ | 0.8596 | 0.9016 | 0.6738 | |
✓ | ✓ | 0.8180 | 0.8323 | 0.6312 | |
✓ | ✓ | ✓ | 0.8644 | 0.9050 | 0.6804 |
dataset | MoE | SRCC | PLCC | KRCC | RMSE |
3k | ✗ | 0.8921 | 0.9257 | 0.7199 | 0.3797 |
✓ | 0.8939 | 0.9273 | 0.7211 | 0.3756 | |
20k | ✗ | 0.8586 | 0.8964 | 0.6712 | 0.3234 |
✓ | 0.8644 | 0.9050 | 0.6804 | 0.3104 |
4.5. Computational Costs
Although integrating LMM features with traditional DNN-based models does introduce additional cost, we emphasize that the incremental cost is minimal: First, the LMM model does not require training and the trained parameters and GFLOPs of ours is similar to those of traditional DNN-based method MANIQA (Parameters: M vs. M, FLOPs: G vs. G). Second, the inference time of extracting LMM features for an image with a resolution of only needs seconds on a NVIDIA RTX 3090 GPU.
4.6. Visualization
To vividly demonstrate the efficacy of the MA-AGIQA framework, we selected 300 images from the AIGCQA-20k and AGIQA-3k datasets where MANIQA had the poorest performance. These images primarily exhibit issues in semantic content. The absolute values of the differences between the model scores and the image ground truth are illustrated in Figure 5, using 0.1 as the bin size for plotting the quality score distribution. The results clearly show that our MA-AGIQA model are more closely aligned with human perception, with a noticeable shift in the difference distribution toward zero and a marked reduction in peak values.
Figure 6 presents a collection of images where the assessments from MANIQA were mostly off the mark. Scores assigned by MANIQA alongside those given by the proposed MA-AGIQA and the ground truth are listed, which reveal that MA-AGIQA markedly enhances alignment with the ground truth in contrast to MANIQA. For instance, in the first image of the top row, MANIQA’s score is 3.50, which diverging substantially from the ground truth score of 1.50. However, MA-AGIQA’s score is 2.98, demonstrating a much closer approximation to the ground truth. This pattern is consistent across all of the images, with MA-AGIQA consistently producing scores that are closer to the ground truth.
5. Conclusion
To mitigate the shortcomings of traditional DNNs in capturing semantic content in AGIs, this study explored the integration of LMMs with traditional DNNs and introduced the MA-AGIQA network. Leveraging mPLUG-Owl2 (Ye et al., 2023), MA-AGIQA efficiently extracts semantic features to enhance quality assessment. The MA-AGIQA network’s ability to dynamically integrate fine-grained semantic features with quality-aware features enables it to effectively handle the varied quality aspects of AGIs. Experiment results across four datasets confirm our model’s superior performance. With thorough ablation studies, the indispensable role of each component within our framework has been validated. This research aspires to catalyze further exploration into the fusion of LMMs within AI-generated content quality assessment and envisions broader application potentials for such methodology.
6. Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 62225112, 62101326, 62301316, and 62271312, the China Postdoctoral Science Foundation under Grants 2023TQ0212 and 2023M742298, and the Postdoctoral Fellowship Program of CPSF under Grant GZC20231618.